# Curriculum Direct Preference Optimization for Diffusion and Consistency Models

### v2

@@ Abstract
Direct Preference Optimization (DPO) has been proposed as an effective and efficient alternative to reinforcement learning from human feedback (RLHF). In this paper, we propose a novel and enhanced version of DPO based on curriculum learning for text-to-image generation. Our method is divided into two training stages. First, a ranking of the examples generated for each prompt is obtained by employing a reward model. Then, increasingly difficult pairs of examples are sampled and provided to a text-to-image generative (diffusion or consistency) model. Generated samples that are far apart in the ranking are considered to form easy pairs, while those that are close in the ranking form hard pairs. In other words, we use the rank difference between samples as a measure of difficulty. The sampled pairs are split into batches according to their difficulty levels, which are gradually used to train the generative model. Our approach, Curriculum DPO, is compared against state-of-the-art fine-tuning approaches on three benchmarks, outperforming the competing methods in terms of text alignment, aesthetics and human preference.
@@ 1 Introduction
Diffusion models [10, 19, 59, 62] represent a family of generative models that gained significant traction in image generation tasks, largely due to their impressive generative capabilities. One of the tasks where these models excel is text-to-image generation [2, 16, 49, 52], as they are capable of generating images that are both aesthetic and well aligned with the input text (prompt). However, state-of-the-art diffusion models that have been widely adopted by the community, e.g. Stable Diffusion [49], are typically heavy in terms of training (and even inference) time. To this end, several studies [6, 14, 35, 70] proposed novel training methods to efficiently fine-tune pre-trained diffusion models. Some of these training frameworks [6, 35, 70] were originally introduced for Large Language Models (LLMs) [8, 21, 46], another family of models that are notoriously hard to train on a few GPUs [57, 68]. This is also the case of Direct Preference Optimization (DPO) [46], a method used to fine-tune LLMs, which was originally proposed as an effective and efficient alternative to reinforcement learning from human feedback (RLHF) [8]. DPO bypasses the need to fit a reward model by harnessing a mapping between reward functions and optimal policies, which enables the direct optimization of the LLM to adhere to human preferences. DPO was later extended to diffusion models [70], showcasing similar benefits in image generation. Despite the significant benefits brought by Diffusion-DPO [70] and related methods [6, 14, 35], there are still observable gaps in terms of various factors, such as text alignment, aesthetics and human preference (see examples in Figures 2, 4 and 5). Therefore, more advanced adaptation techniques are required to reduce these gaps.
@@
We carry out experiments across three evaluation benchmarks to assess the effectiveness of Curriculum DPO in adapting text-to-image generative models to different reward models. We first tackle the text alignment task, employing the Sentence-BERT [48] similarity metric to evaluate the correspondence between the actual text prompt and the caption provided by the LLaVA model [31] for each generated image. Next, we focus on enhancing the visual aesthetics of the images, where we use the LAION Aesthetic Predictor [55] as the reward model. Finally, for the third task, we increase the human preference rate by employing HPSv2 [73] as the reward model. The quantitative and qualitative results on these three benchmarks show that Curriculum DPO surpasses state-of-the-art fine-tuning strategies, such as Diffusion-DPO [46, 70] and Denoising Diffusion Policy Optimization (DDPO) [6]. Furthermore, we conduct a study based on actual human feedback to further evaluate the fine-tuning methods. The outcome of this study reinforces the findings from our quantitative analysis, showing that Curriculum DPO consistently generates the most preferred samples out of the three methods. Overall, our results demonstrate the effectiveness of Curriculum DPO. By introducing increasingly complex pairs, Curriculum DPO ensures that the generative models progressively refine their capabilities, leading to outputs that are better aligned with the reward models.
@@
  • We demonstrate the superiority of Curriculum DPO over state-of-the-art training alternatives on three evaluation benchmarks.
@@ 2 Related Work
Recent breakthroughs in LLMs were adopted and applied to diffusion models. Hu et al. [21] proposed an efficient method to fine-tune LLMs by introducing a trainable low-rank matrix decomposition for certain parameters of transformer layers, and then summing them, while keeping the original weights frozen. Luo et al. [35] employed Low-Rank Adaptation (LoRA) [21] on consistency models, demonstrating better results while requiring less memory. The loss function for training LLMs is not capturing human preference, while current metrics (e.g. BLEU [42]) defectively assess it. To this end, reinforcement learning methods have been adopted [3, 17, 77]. By learning a reward function from a dataset collected with human feedback and fine-tuning a text-to-image diffusion model on it, Lee et al. [29] showed that the generated images are better aligned with the input prompt. Fine-tuning LLMs using policy gradient algorithms has played a major role in their success, and thus, they were rapidly adopted in diffusion models as well. While Fan et al. [14] demonstrated that such an algorithm leads to an improved text-image alignment, Black et al. [6] applied their method on enhancing multiple aspects of synthesis: compressibility, aesthetics, as well as prompt alignment. In their recent work, Rafailov et al. [46] presented Direct Preference Optimization, an algorithm for fine-tuning LLMs from any preference, proving to be superior to Proximal Policy Optimization methods [41, 56, 77], and thus, replacing RLHF algorithms. It involves directly training the model by modifying the objective to integrate the preference rather than first fitting a reward model and then using it to train the original model. Wallace et al. [70] reformulated DPO to fit the context of diffusion models by employing a novel training strategy and loss function. Their experiments demonstrate promising results on subjective metrics such as visual aesthetics or prompt alignment. Nevertheless, to the best of our knowledge, curriculum learning has not been applied in conjunction with controllable generation methods, such as DPO [70] and DDPO [6].
@@ 3 Method
We begin by introducing the mathematical preliminaries behind diffusion models, consistency models, DPO and Diffusion-DPO (for a more detailed introduction of these concepts, please refer to Appendix A.1). We then present our adaptation of DPO for consistency models. Finally, we delve into our curriculum learning strategy designed for DPO.
@@ 3.1 Preliminaries
$\mathcal{L}_{\scriptsize{\mbox{simple}}}=\mathbb{E}_{t\sim\mathcal{U}(1,T),\epsilon\sim\mathcal{N}(0,\mathbf{I}),x_{0}\sim p(x_{0})}\!\left\lVert\epsilon_{t}-\epsilon_{\theta}(x_{t},t)\right\rVert^{2}_{2},$ (1)
where $\epsilon_{\theta}(x_{t},t)$ represents the time-dependent neural network and $\theta$ are the trainable parameters.
The generation of new samples with diffusion models is a multi-step process where an initial Gaussian sample is progressively transformed into one from the original distribution. At each step, the model $\epsilon_{\theta}$ estimates and subtracts the noise, refining the sample towards the target with each iteration.
@@
$\mathcal{L}_{\scriptsize{\mbox{CD}}}(\phi)=d(f_{\phi}(x_{t_{n+1}},t_{n+1}),f_{\phi^{-}}(\hat{x}_{t_{n}}^{\theta},t_{n})),$ (2)
@@
Direct Preference Optimization (DPO). Rafailov et al. [46] introduced DPO, a method for fine-tuning LLMs with pairs of ranked examples $(x_{0}^{w},x_{0}^{l},c)$, where $x_{0}^{w}$ is preferred, $x_{0}^{l}$ is less favored, and $c$ is a condition under which both samples are generated. The training objective aims to increase the likelihood of the preferred examples, $p_{\theta}(x_{0}^{w}|c)$, and decrease that of less preferred ones, while not diverging from the initial state of the model, denoted by $p_{\scriptsize{\mbox{ref}}}(x_{0}|c)$:
$\mathcal{L}_{\mathrm{\scriptsize{\mbox{DPO}}}}(\theta)=-\mathbb{E}_{x_{0}^{w},x_{0}^{l},c}\left(\log{\sigma\left(\beta\cdot\left(\log{\frac{p_{\theta}(x_{0}^{w}|c)}{p_{\scriptsize{\mbox{ref}}}(x_{0}^{w}|c)}}-\log{\frac{p_{\theta}(x_{0}^{l}|c)}{p_{\scriptsize{\mbox{ref}}}(x_{0}^{l}|c)}}\right)\right)}\right),$ (3)
where $\sigma$ is the sigmoid function and $\beta$ is a hyperparameter controlling the divergence of $p_{\theta}(x_{0}|c)$ from $p_{\scriptsize{\mbox{ref}}}(x_{0}|c)$. To better grasp the intuition behind $\mathcal{L}_{\mathrm{\scriptsize{\mbox{DPO}}}}$, we can look at its gradient with respect to $\theta$:
$\frac{\partial\mathcal{L}_{\mathrm{\scriptsize{\mbox{DPO}}}}(\theta)}{\partial\theta}=-\beta\mathbb{E}_{x_{0}^{w},x_{0}^{l},c}\left[\sigma\left(\hat{r}_{\theta}(x_{0}^{l},c)-\hat{r}_{\theta}(x_{0}^{w},c)\right)\cdot\left(\frac{\partial\log{p_{\theta}(x_{0}^{w}|c)}}{\partial\theta}-\frac{\partial\log{p_{\theta}(x_{0}^{l}|c)}}{\partial\theta}\right)\right],$ (4)
where $\hat{r}_{\theta}(x_{0},c)=\beta\cdot\log{\frac{p_{\theta}(x_{0}|c)}{p_{\tiny{\mbox{ref}}}(x_{0}|c)}}$. By analyzing Eq. (4), we can observe that the DPO objective enhances the likelihood of preferred examples while diminishing it for the unfavored ones. Moreover, the update is weighted by a sigmoid term which is proportional to how wrong the model is.
@@
$\begin{split}\mathcal{L}_{\mathrm{\scriptsize{\mbox{Diff-DPO}}}}(\theta)=-\mathbb{E}_{x_{t}^{w},x_{t}^{l},c}\Big{[}\log\sigma\Big{(}\!-\beta T\Big{(}&\left(\lVert\epsilon^{w}-\epsilon_{\theta}^{w}(x_{t}^{w},t,c)\rVert_{2}^{2}-\lVert\epsilon^{w}-\epsilon_{ref}^{w}(x_{t}^{w},t,c)\rVert_{2}^{2}\right)-\\ &\left(\lVert\epsilon^{l}-\epsilon_{\theta}^{l}(x_{t}^{l},t,c)\rVert_{2}^{2}-\lVert\epsilon^{l}-\epsilon_{ref}^{l}(x_{t}^{l},t,c)\rVert_{2}^{2}\right)\Big{)}\Big{)}\Big{]},\end{split}$ (5)
@@ 3.2 Consistency-DPO
Inspired by the DPO implementation of Wallace et al. [70] for diffusion models, we propose a parallel adaptation for consistency models. Specifically, we utilize Eq. (2) to measure and enhance the model’s accuracy on preferred examples, while permitting a rise in error rates for less favored ones. Given a reference pre-trained consistency model $f_{\scriptsize{\mbox{ref}}}$, we propose the following DPO objective for fine-tuning on ranked examples:
$\mathcal{L}_{\mathrm{\scriptsize{\mbox{Con-DPO}}}}(\phi)=-\mathbb{E}_{x_{t_{n+1}}^{w},x_{t_{n+1}}^{l},c}\Big{[}\log\sigma\Big{(}-\beta(d^{w}(x_{t_{n+1}}^{w},\hat{x}_{t_{n}}^{w,\theta},\phi)-d^{l}(x_{t_{n+1}}^{l},\hat{x}_{t_{n}}^{l,\theta},\phi))\Big{)}\Big{]},$ (6)
@@
$d^{*}=d(f_{\phi}(x_{t_{n+1}}^{*},t_{n+1},c),f_{\scriptsize{\mbox{ref}}}(\hat{x}_{t_{n}}^{*,\theta},t_{n},c))-d(f_{\scriptsize{\mbox{ref}}}(x_{t_{n+1}}^{*},t_{n+1},c),f_{\scriptsize{\mbox{ref}}}(\hat{x}_{t_{n}}^{*,\theta},t_{n},c)).$ (7)
Note that $c$ is a condition, as in previous cases. The variable $x_{t_{n+1}}^{*}$ is derived using the forward process of a diffusion model, applied on $x_{0}^{*}$. Additionally, $\hat{x}_{t_{n}}^{*,\theta}$ is the result of one discretization step of an ODE solver applied to the PF-ODE , starting from $x_{t_{n+1}}^{*}$ and using a pre-trained diffusion model parameterized by $\theta$. Moreover, $d(\cdot,\cdot)$ is a distance measure, $\sigma$ is the sigmoid function, $n\sim\mathcal{U}(1,N)$, $N$ is the discretization length of the interval $[0,T]$, and $\phi$ are the trainable parameters of the consistency model. We did not use a running average of $\phi$ for the Consistency-DPO loss because we already have the pre-trained model $f_{\scriptsize{\mbox{ref}}}$, which has the necessary self-consistency property.
To show that the Consistency-DPO loss defined in Eq. (6) has similar properties to the original DPO and Diffusion-DPO losses defined in Eq. (3) and Eq. (5), respectively, we can consider the gradient with respect to model’s parameters:
$\frac{\partial\mathcal{L}_{\mathrm{\scriptsize{\mbox{Con-DPO}}}}}{\partial\phi}=\beta\mathbb{E}_{x_{t_{n+1}}^{w},x_{t_{n+1}}^{l},c}\left[\sigma\left(\beta\left(d^{w}-d^{l}\right)\right)\left(\frac{\partial d^{w}}{\partial\phi}-\frac{\partial d^{l}}{\partial\phi}\right)\right].$ (8)
The gradient described in Eq. (8) exhibits similar properties to the gradient outlined in Eq. (4). Specifically, it reduces the distance metric $d$ between $f_{\phi}$ and $f_{\scriptsize{\mbox{ref}}}$ for the favored examples and, at the same time, it increases the same metric for less preferred examples. Additionally, this gradient is also weighted by the sigmoid function, which is close to $1$ when $f_{\phi}$ shows a weaker preference for the favored example ($x_{t_{n+1}}^{w}$) than the level of preference yielded by $f_{\scriptsize{\mbox{ref}}}$.
3.3 Curriculum DPO
@@
We formally describe Curriculum DPO in Algorithm 1. In step 1, the set of examples $\{x_{0,i}\}_{i=1}^{M}$ that are generated under a condition $c$ (i.e. a text prompt) are sorted based on the preferences indicated by a reward model $r_{\varphi}$, such that the ranking follows $r_{\varphi}(x_{0,1},c)\geq\dots\geq r_{\varphi}(x_{0,M},c)$. In a typical DPO setup, the pairs are formed from the set defined in step 2 of the algorithm, $S=\left\{(x_{0,i},x_{0,j},c)|i,j\in\left\{1,\dots,M\right\},i<j,r_{\varphi}(x_{0,i},c)>r_{\varphi}(x_{0,j}^{l},c)\right\}$. In steps 3-5, Curriculum DPO segments this set into $B$ distinct batches, where each batch is denoted by $S_{k}=\left\{(x_{0}^{w},x_{0}^{l},c)|(x_{0}^{w},x_{0}^{l})=(x_{0,i},x_{0,j});\mathrm{L}_{k}<j-i\leq\mathrm{R}_{k};(x_{0,i},x_{0,j},c)\in S\right\}$, where $k=\{1,2,...,B\}$. Each batch $S_{k}$ (not to be confused with mini-batches) is defined by the limits $\mathrm{L}_{k}$ and $\mathrm{R}_{k}$, which represent the smallest and largest possible differences between positions $j$ and $i$ within the ranking made by the reward model. The training is executed in steps 7-16. At each iteration, the current batch $S_{k}$ is first included in the training set $P$. When $k=1$, the training set will comprise the first batch, $S_{1}$, which will be used to gradually adapt the generative model to the easiest (most distinctive) pairs, enhancing the training efficiency during the early learning iterations. After adding the current batch $S_{k}$ to the training set at step 8, the generative model is trained for $H_{k}$ iterations on $P$. The training iterations (steps 10-16) are custom to each specific DPO implementation, namely Diffusion-DPO and Consistency-DPO. In Algorithm 1, we present the iterations that are specific to the Consistency-DPO approach. In Appendix A.2, we present the analogous version for Diffusion-DPO.

### v3

@@ Abstract
Direct Preference Optimization (DPO) has been proposed as an effective and efficient alternative to reinforcement learning from human feedback (RLHF). In this paper, we propose a novel and enhanced version of DPO based on curriculum learning for text-to-image generation. Our method is divided into two training stages. First, a ranking of the examples generated for each prompt is obtained by employing a reward model. Then, increasingly difficult pairs of examples are sampled and provided to a text-to-image generative (diffusion or consistency) model. Generated samples that are far apart in the ranking are considered to form easy pairs, while those that are close in the ranking form hard pairs. In other words, we use the rank difference between samples as a measure of difficulty. The sampled pairs are split into batches according to their difficulty levels, which are gradually used to train the generative model. Our approach, Curriculum DPO, is compared against state-of-the-art fine-tuning approaches on nine benchmarks, outperforming the competing methods in terms of text alignment, aesthetics and human preference. Our code is available at https://github.com/CroitoruAlin/Curriculum-DPO.
@@ 1 Introduction
Diffusion models [11, 20, 61, 64] represent a family of generative models that gained significant traction in image generation tasks, largely due to their impressive generative capabilities. One of the tasks where these models excel is text-to-image generation [3, 17, 51, 54], as they are capable of generating images that are both aesthetic and well aligned with the input text (prompt). However, state-of-the-art diffusion models that have been widely adopted by the community, e.g. Stable Diffusion [51], are typically heavy in terms of training (and even inference) time. To this end, several studies [7, 15, 37, 72] proposed novel training methods to efficiently fine-tune pre-trained diffusion models. Some of these training frameworks [7, 37, 72] were originally introduced for Large Language Models (LLMs) [9, 22, 48], another family of models that are notoriously hard to train on a few GPUs [59, 70]. This is also the case of Direct Preference Optimization (DPO) [48], a method used to fine-tune LLMs, which was originally proposed as an effective and efficient alternative to reinforcement learning from human feedback (RLHF) [9]. DPO bypasses the need to fit a reward model by harnessing a mapping between reward functions and optimal policies, which enables the direct optimization of the LLM to adhere to human preferences. DPO was later extended to diffusion models [72], showcasing similar benefits in image generation. Despite the significant benefits brought by Diffusion-DPO [72] and related methods [7, 15, 37], there are still observable gaps in terms of various factors, such as text alignment, aesthetics and human preference (see Figures 2, 5 and 6). Therefore, more advanced adaptation techniques are required to reduce these gaps.
@@
We carry out experiments across nine evaluation benchmarks to assess the effectiveness of Curriculum DPO in adapting text-to-image generative models to different reward models. We first tackle the text alignment task, employing the Sentence-BERT [50] similarity metric to evaluate the correspondence between the actual text prompt and the caption provided by the LLaVA model [33] for each generated image. Next, we focus on enhancing the visual aesthetics of the images, where we use the LAION Aesthetics Predictor [57] as the reward model. Finally, for the third task, we increase the human preference rate by employing HPSv2 [75] as the reward model. Each of the three tasks is studied on two distinct datasets. The quantitative and qualitative results on all datasets show that Curriculum DPO surpasses state-of-the-art fine-tuning strategies, such as Diffusion-DPO [48, 72] and Denoising Diffusion Policy Optimization (DDPO) [7]. Furthermore, we conduct a study based on actual human feedback to further evaluate the fine-tuning methods. The outcome of this study reinforces the findings from our quantitative analysis, showing that Curriculum DPO consistently generates the most preferred samples out of the three methods. Moreover, our ablation study on the number of training samples shows that Curriculum DPO achieves similar performance to Diffusion-DPO and DDPO, while using 10$\times$ less images. Overall, our results demonstrate the effectiveness of Curriculum DPO. By introducing increasingly complex pairs, Curriculum DPO ensures that the generative models progressively refine their capabilities, leading to outputs that are better aligned with the reward models.
@@
  • We demonstrate the superiority of Curriculum DPO over state-of-the-art training alternatives on nine evaluation benchmarks.
@@ 2 Related Work
Recent breakthroughs in LLMs were adopted and applied to diffusion models. Hu et al. [22] proposed an efficient method to fine-tune LLMs by introducing a trainable low-rank matrix decomposition for certain parameters of transformer layers, and then summing them, while keeping the original weights frozen. Luo et al. [37] employed Low-Rank Adaptation (LoRA) [22] on consistency models, demonstrating better results while requiring less memory. The loss function for training LLMs is not capturing human preference, while current metrics (e.g. BLEU [44]) defectively assess it. To this end, reinforcement learning methods have been adopted [4, 18, 79]. By learning a reward function from a dataset collected with human feedback and fine-tuning a text-to-image diffusion model on it, Lee et al. [31] showed that the generated images are better aligned with the input prompt. Fine-tuning LLMs using policy gradient algorithms has played a major role in their success, and thus, they were rapidly adopted in diffusion models as well. While Fan et al. [15] demonstrated that such an algorithm leads to an improved text-image alignment, Black et al. [7] applied their method on enhancing multiple aspects of synthesis: compressibility, aesthetics, as well as prompt alignment. In their recent work, Rafailov et al. [48] presented Direct Preference Optimization, an algorithm for fine-tuning LLMs from any preference, proving to be superior to Proximal Policy Optimization methods [43, 58, 79], and thus, replacing RLHF algorithms. It involves directly training the model by modifying the objective to integrate the preference rather than first fitting a reward model and then using it to train the original model. Wallace et al. [72] reformulated DPO to fit the context of diffusion models by employing a novel training strategy and loss function. Their experiments demonstrate promising results on subjective metrics such as visual aesthetics or prompt alignment. Nevertheless, to the best of our knowledge, curriculum learning has not been applied in conjunction with controllable generation methods for diffusion models, such as DPO [72] and DDPO [7].
@@ 3 Method
We begin by introducing the mathematical preliminaries behind diffusion models, consistency models, DPO and Diffusion-DPO (for a more detailed introduction of these concepts, please refer to Appendix 6). We then present our adaptation of DPO for consistency models. Finally, we delve into our curriculum learning strategy designed for DPO.
@@ 3.1 Preliminaries
$\!\!\mathcal{L}_{\mbox{\scriptsize{simple}}}\!=\!\mathbb{E}_{t\sim\mathcal{U}(1,T),\epsilon\sim\mathcal{N}(0,\mathbf{I}),x_{0}\sim p(x_{0})}\!\left\lVert\epsilon_{t}\!-\!\epsilon_{\theta}(x_{t},t)\right\rVert^{2}_{2}\!,$ (1)
where $\epsilon_{\theta}(x_{t},t)$ represents the time-dependent neural network and $\theta$ are the trainable parameters. The generation of new samples with diffusion models is a multi-step process where an initial Gaussian sample is progressively transformed into one from the original distribution. At each step, the model $\epsilon_{\theta}$ estimates and subtracts the noise, refining the sample towards the target with each iteration.
@@
$\mathcal{L}_{\mbox{\scriptsize{CD}}}(\phi)=d(f_{\phi}(x_{t_{n+1}},t_{n+1}),f_{\phi^{-}}(\hat{x}_{t_{n}}^{\theta},t_{n})),$ (2)
@@
Direct Preference Optimization (DPO). Rafailov et al. [48] introduced DPO, a method to fine-tune LLMs with pairs of ranked examples $(x_{0}^{w},x_{0}^{l},c)$, where $x_{0}^{w}$ is preferred, $x_{0}^{l}$ is less favored, and $c$ is a condition under which both samples are generated. The training objective aims to increase the likelihood of the preferred examples, $p_{\theta}(x_{0}^{w}|c)$, and decrease that of less preferred ones, while not diverging from the initial state of the model, denoted by $p_{\mbox{\scriptsize{ref}}}(x_{0}|c)$:
$\begin{split}\mathcal{L}_{\mathrm{\mbox{\scriptsize{DPO}}}}(\theta)\!\!=\!\!-\!\mathbb{E}_{x_{0}^{w}\!,x_{0}^{l},c}\!\left[\!\log\!\sigma\!\bigg{(}\!\!\beta\!\!\left(\!\!\log\!{\frac{p_{\theta}(x_{0}^{w}\!|c)}{p_{\mbox{\scriptsize{ref}}}(x_{0}^{w}\!|c)}}\!-\!\log\!{\frac{p_{\theta}(x_{0}^{l}\!|c)}{p_{\mbox{\scriptsize{ref}}}(x_{0}^{l}\!|c)}}\!\!\right)\!\!\!\!\bigg{)}\!\!\right]\!\!,\end{split}$ (3)
where $\sigma$ is the sigmoid function and $\beta$ is a hyperparameter controlling the divergence of $p_{\theta}(x_{0}|c)$ from $p_{\mbox{\scriptsize{ref}}}(x_{0}|c)$. To better grasp the intuition behind $\mathcal{L}_{\mathrm{\mbox{\scriptsize{DPO}}}}$, we can look at its gradient with respect to $\theta$:
$\begin{split}\frac{\partial\mathcal{L}_{\mathrm{\mbox{\scriptsize{DPO}}}}(\theta)}{\partial\theta}\!&=\!-\beta\mathbb{E}_{x_{0}^{w},x_{0}^{l},c}\Bigg{[}\sigma\left(\hat{r}_{\theta}(x_{0}^{l},c)-\hat{r}_{\theta}(x_{0}^{w},c)\right)\cdot\\ &\left(\frac{\partial\log{p_{\theta}(x_{0}^{w}|c)}}{\partial\theta}-\frac{\partial\log{p_{\theta}(x_{0}^{l}|c)}}{\partial\theta}\right)\!\Bigg{]},\end{split}$ (4)
where $\hat{r}_{\theta}(x_{0},c)=\beta\cdot\log{\frac{p_{\theta}(x_{0}|c)}{p_{\mbox{\tiny{ref}}}(x_{0}|c)}}$. By analyzing Eq. (4), we can observe that the DPO objective enhances the likelihood of preferred examples while diminishing it for the unfavored ones. Moreover, the update is weighted by a sigmoid term that is proportional to how wrong the model is.
@@
$\begin{split}\!\mathcal{L}_{\mathrm{\mbox{\scriptsize{Diff-DPO}}}}&(\theta)\!=\!-\mathbb{E}_{x_{t}^{w},x_{t}^{l},c}\!\Big{[}\!\log\sigma\Big{(}\!\!-\!\beta\cdot T\Big{(}\\ &\lVert\epsilon^{w}\!-\!\epsilon_{\theta}^{w}(x_{t}^{w}\!,t,c)\rVert_{2}^{2}\!-\!\lVert\epsilon^{w}\!-\!\epsilon_{\mbox{\scriptsize{ref}}}^{w}(x_{t}^{w},t,c)\rVert_{2}^{2}-\\ &\lVert\epsilon^{l}\!-\!\epsilon_{\theta}^{l}(x_{t}^{l},t,c)\rVert_{2}^{2}\!+\!\lVert\epsilon^{l}\!-\!\epsilon_{\mbox{\scriptsize{ref}}}^{l}(x_{t}^{l},t,c)\rVert_{2}^{2}\!\Big{)}\!\Big{)}\!\Big{]}\!,\end{split}$ (5)
@@ 3.2 Consistency-DPO
Inspired by the DPO implementation of Wallace et al. [72] for diffusion models, we propose a parallel adaptation for consistency models. Specifically, we utilize Eq. (2) to measure and enhance the model’s accuracy on preferred examples, while permitting a rise in error rates for less favored ones. Given a reference pre-trained consistency model $f_{\mbox{\scriptsize{ref}}}$, we propose the following DPO objective to fine-tune on ranked examples:
$\begin{split}\mathcal{L}_{\mathrm{\mbox{\scriptsize{Con-DPO}}}}&(\phi)\!=\!-\mathbb{E}_{x_{t_{n+1}}^{w},x_{t_{n+1}}^{l},c}\Big{[}\log\sigma\Big{(}\!-\beta\Big{(}\\ &d^{w}(x_{t_{n+1}}^{w},\hat{x}_{t_{n}}^{w,\theta},\phi)-d^{l}(x_{t_{n+1}}^{l},\hat{x}_{t_{n}}^{l,\theta},\phi)\!\Big{)}\!\Big{)}\!\Big{]},\end{split}$ (6)
@@
$\begin{split}d^{*}=&\,d(f_{\phi}(x_{t_{n+1}}^{*},t_{n+1},c),f_{\mbox{\scriptsize{ref}}}(\hat{x}_{t_{n}}^{*,\theta},t_{n},c))-\\ &\,d(f_{\mbox{\scriptsize{ref}}}(x_{t_{n+1}}^{*},t_{n+1},c),f_{\mbox{\scriptsize{ref}}}(\hat{x}_{t_{n}}^{*,\theta},t_{n},c)).\end{split}$ (7)
Note that $c$ is a condition, as in previous cases. The variable $x_{t_{n+1}}^{*}$ is derived using the forward process of a diffusion model, applied to $x_{0}^{*}$. Additionally, $\hat{x}_{t_{n}}^{*,\theta}$ is the result of one discretization step of an ODE solver applied to the PF-ODE , starting from $x_{t_{n+1}}^{*}$ and using a pre-trained diffusion model parameterized by $\theta$. Moreover, $d(\cdot,\cdot)$ is a distance measure, $\sigma$ is the sigmoid function, $n\sim\mathcal{U}(1,N)$, $N$ is the discretization length of the interval $[0,T]$, and $\phi$ are the trainable parameters of the consistency model. We did not use a running average of $\phi$ for the Consistency-DPO loss because we already have the pre-trained model $f_{\mbox{\scriptsize{ref}}}$, which has the necessary self-consistency property.
@@
To show that the Consistency-DPO loss defined in Eq. (6) has similar properties to the original DPO and Diffusion-DPO losses defined in Eq. (3) and Eq. (5), respectively, we can consider the gradient with respect to $\phi$:
$\begin{split}\frac{\partial\mathcal{L}_{\mathrm{\mbox{\scriptsize{Con-DPO}}}}}{\partial\phi}\!=\!\beta\mathbb{E}_{x_{t_{n+1}}^{w}\!,x_{t_{n+1}}^{l}\!,c}\!\Bigg{[}\!\sigma\!\left(\beta\!\left(d^{w}\!-\!d^{l}\right)\!\right)\!\!\left(\!\frac{\partial d^{w}}{\partial\phi}\!-\!\frac{\partial d^{l}}{\partial\phi}\!\right)\!\!\!\Bigg{]}\!.\end{split}$ (8)
The gradient described in Eq. (8) exhibits similar properties to the gradient outlined in Eq. (4). Specifically, it reduces the distance metric $d$ between $f_{\phi}$ and $f_{\mbox{\scriptsize{ref}}}$ for the favored examples and, at the same time, it increases the same metric for less preferred examples. Additionally, this gradient is also weighted by the sigmoid function, which is close to $1$ when $f_{\phi}$ shows a weaker preference for the favored example ($x_{t_{n+1}}^{w}$) than the level of preference yielded by $f_{\mbox{\scriptsize{ref}}}$.
3.3 Curriculum DPO
We formally describe Curriculum DPO for consistency models in Algorithm 1. In step 1, the set of examples $\{x_{0,i}\}_{i=1}^{M}$ that are generated under a condition $c$ (i.e. a text prompt) are sorted based on the preferences indicated by a reward model $r_{\varphi}$, such that the ranking follows $r_{\varphi}(x_{0,1},c)\geq\dots\geq r_{\varphi}(x_{0,M},c)$. In a typical DPO setup, the pairs are formed from the set defined in step 2 of the algorithm, denoted by $S$. In steps 3-5, Curriculum DPO segments this set into $B$ distinct batches, where each batch is denoted by $S_{k}$, where $k=\{1,2,...,B\}$. Each batch $S_{k}$ (not to be confused with mini-batches) is defined by the limits $\mathrm{L}_{k}$ and $\mathrm{R}_{k}$, which represent the smallest and largest possible differences between positions $j$ and $i$ within the ranking made by the reward model. The training is executed in steps 7-16. At each iteration, the current batch $S_{k}$ is first included in the training set $P$. When $k=1$, the training set will comprise the first batch, $S_{1}$, which will be used to gradually adapt the generative model to the easiest (most distinctive) pairs, enhancing the training efficiency during the early learning iterations. After adding the current batch $S_{k}$ to the training set at step 8, the generative model is trained for $H_{k}$ iterations on $P$. The training iterations (steps 10-16) are custom to each specific DPO implementation, namely Diffusion-DPO and Consistency-DPO. In Algorithm 1, we present the iterations that are specific to the Consistency-DPO approach. In Appendix 7, we present the analogous version for Diffusion-DPO.